determine the equation of line passing through the point (-6,-2) that is perpendicular to y=2x-4
Answer by barunjha(4)

Ans:- x+2y+10=0
Explanation-
Given equation of line is y=2x-4.
comparing this line with standard equation of line in slope intercept form
y=mx+c.
here m= 2.
let the equation of required line is y=Mx+k.

from condition of perpendicularity of two straight line m*M= -1.
M= -1/m. i.e -1/2.
putting the value of M in required equation.
y=-x/2+k
i.e 2y= -x+K, where 2k= K.
as this passes through the point ( -6,-2) so this will satisfy the equation.
hence 2(-2)= -(-6)+K.
solving it for K= 10
so Equation for new line will be
x+2y+10=0.